OpenStudy (anonymous):

Find the first 5 terms of this series which centers at x=0

OpenStudy (anonymous):



OpenStudy (anonymous):

Summation

OpenStudy (anonymous):

I'm thinking of starting with sin(x) = the series of (-1)^n * x^(2n+1) / (2n+1)!

OpenStudy (amistre64):

thats a good start yes

OpenStudy (amistre64):

the 1/5 is a constant so we can set it aside

OpenStudy (anonymous):

Okay and then sin(x) / 5x would be 1/2 * the summation of (-1)^n * x^(2n) / (2n+1)!

OpenStudy (anonymous):

1/5*

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Okay, so then you just plug in n=1,2,3,4,5...?

OpenStudy (amistre64):

n=0,1,2,3,4 is what id do, the summation of sin starts at n=0 in your presentation there

OpenStudy (anonymous):

Wait but isn't an integral? So we have to find the integral of that summation?

OpenStudy (amistre64):

thats what you posted .... no take backs lol

OpenStudy (anonymous):

Ah no, I was just catching my mistake xD

OpenStudy (anonymous):

Okay how would you find the integral of this:

OpenStudy (anonymous):

\[\sum_{}^{} (-1)^n * x^{2n} / (2n+1)!\]

OpenStudy (amistre64):

you have a polynomial, how do you integrate powers of x?

OpenStudy (amistre64):

all the other stuff is just constants

OpenStudy (anonymous):

Ah you just add +1 to the powers of x...

OpenStudy (amistre64):

and divide, yeah

OpenStudy (anonymous):

\[\sum_{}^{} (-1)^n * x^{2n+1} / (2n+1)! + c\]

OpenStudy (anonymous):

Okay but wolfram alpha states that there is a (2n+1) on the bottom.... how?

OpenStudy (amistre64):

for starter, you didnt divide your exponent x^n integrates to x^(n+1)/(n+1)

OpenStudy (amistre64):

and there is no +C in this case, unless you want to define it as the constant first term.

OpenStudy (anonymous):

Oh I get it! thanks :)

OpenStudy (amistre64):

\[\int\frac{(-1)^n}{(2n+1)!}x^{2n}dx\] the only thing that is affected by x, is x itself \[\frac{(-1)^n}{(2n+1)(2n+1)!}x^{2n+1}\] \[\frac{(-1)^n}{(2n+1)(2n+1)(2n)!}x^{2n+1}\] \[\frac{(-1)^n}{(2n+1)^2(2n)!}x^{2n+1}\] but thats just playing with algebra

OpenStudy (amistre64):

if we simply list the first 5 terms \[\frac{1}{1!},\frac{-1}{3!}x^2,\frac{1}{5!}x^4,\frac{-1}{7!}x^6,\frac{-1}{9!}x^8\] integrate them \[\frac{1}{1!}x,\frac{-1}{3.3!}x^3,\frac{1}{5.5!}x^5,\frac{-1}{7.7!}x^7,\frac{-1}{9.9!}x^9\]

OpenStudy (amistre64):

lol, got a stray negative in there

OpenStudy (amistre64):

dont forget the 1/5 laying outside